OpenStudy (anonymous):

what is the reference angle of -775? a) 550° b) -305° c) -55° d) 305°

OpenStudy (anonymous):

d)

OpenStudy (anonymous):

its c)

OpenStudy (anonymous):

mathod for any angle lower than 0 we add 360 till it gets between 0 and 360

OpenStudy (anonymous):

can you show the work?

OpenStudy (anonymous):

sorry - i got it wrong - i'd forgotten this bit of trigonometry. i.ve just checked it out: the reference angle is the acute angle formed with reference to x axis. for example reference angle for 120 degress is the acute angles in second quadrant which is 180-120 = 60 degrees. if the angle is greater than 360 then you subtract 360's until you get a value between 0 and 360 eg 380 - refernce angle is 380-36 = 20 degrees. for 480 it is 480-360 = 120 then 180 - 120 = 60 degrees. 790 - 360 - 360 = 70 degrees.

OpenStudy (anonymous):

refer to www.analyzemath.com/Calculators/Reference_Cal.html for negative angles this site takes the positive acute angle and returns 55 degrees for -775

OpenStudy (anonymous):

not -55!